# A Comprehensive Study on Large Language Models for Mutation Testing

### v4

@@ 4. Discussion
In this section, we explore the impact of different experiment setups for LLMs, then we discuss the implications.
@@ 4.1.1. Setup of the Prompts
Prompts play a critical role in the performance of LLMs. Therefore, we aim to investigate how different prompts influence mutation generation effectiveness. Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
For the prompts exploration, we use two representative LLMs, i.e., the closed-source GPT-3.5-Turbo and the recent open-source model DeepSeek-V3-761b, then switch different prompt templates. More specifically, we modify the default prompt template by adding and removing different sources of information in the Context, and we finally get five types of prompts, shown as follows:
Prompt 1 (P1): The Default Prompt. P1 is the default prompt, shown as Figure 1.
Prompt 2 (P2): P1 without Few-Shot Examples. P2 is based on P1 by removing all the few-shot examples.
Prompt 3 (P3): P2 without Whole Java Method. P3 is based on P2 by removing the surrounding Java method code, only keeping the target code element to be mutated.
Prompt 4 (P4): P1 with Unit Tests. Based on the default prompt P1, P4 further adds the source code of the corresponding unit tests of the target method into Context.
Prompt 5 (P5): P1 with Whole Java File. P5 is a variant of P1, which uses the code of the entire Java file as the code context, rather than the entire method.
| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||||||
@@
| Mut. # | 3681 | 3688 | 3350 | 3210 | 3496 | 3851 | 3847 | 3834 | 3844 | 3878 |
| Mut. Score | 0.773 | 0.785 | 0.842 | 0.801 | 0.694 | 0.765 | 0.750 | 0.760 | 0.753 | 0.776 |
| Real Bug Detec. | 79.5% | 78.4% | 72.2% | 72.5% | 79.5% | 85.5% | 85.8% | 83.9% | 84.7% | 83.1% |
| Avg. Ochiai | 63.1% | 61.8% | 60.8% | 62.1% | 62.3% | 67.5% | 65.6% | 66.4% | 65.8% | 66.4% |
| Coupling Rate | 52.1% | 52.2% | 47.4% | 49.6% | 47.3% | 53.2% | 51.7% | 50.9% | 50.0% | 49.1% |
| Compling Rate | 66.9% | 66.8% | 60.0% | 61.2% | 70.3% | 82.8% | 83.5% | 78.0% | 82.0% | 79.5% |
| Dup. Mut. Rate | 11.4% | 13.3% | 14.9% | 12.7% | 16.9% | 7.3% | 6.4% | 8.5% | 7.0% | 8.2% |
| Avg. Gen. Time | 1.789 | 1.782 | 1.09 | 1.884 | 1.907 | 2.573 | 2.571 | 2.488 | 2.601 | 2.596 |
| Avg. # Token | 105 | 95 | 44 | 219 | 393 | 104 | 92 | 43 | 219 | 392 |
@@
Note: The prompt template that achieves the best performance is highlighted in grey background color.
Table 14 presents a comparative evaluation of different prompts for GPT-3.5 and DeepSeek-V3-761b across multiple metrics. Across both models, P1 delivers the best overall performance on all effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), demonstrating its strength in generating mutations that are semantically closer to real bugs. For the validity metrics, P1 ranks second in compilation rate on GPT-3.5 and first on DeepSeek-V3-761b, while achieving the lowest duplicate mutation rate on GPT-3.5 and third-lowest on DeepSeek-V3-761b. Notably, compared with P5, which provides the entire file as context, P1 relies only on the full method yet achieves higher semantic similarity to real bugs, with reduced generation time and token usage, while maintaining comparable compilation and duplicate mutation rates.
Across P1 to P5, we observe a clear context–performance trade-off. Removing relevant context (P3) substantially degrades effectiveness, as seen in lower real bug detectability, average Ochiai, and coupling rates. Providing a focused, method-level context (P1 and P2) delivers the best overall balance of effectiveness and validity. However, adding excessive context beyond the focal method (P4 and P5) does not yield further improvements and often increases token usage and generation time, while in some cases even reducing key effectiveness metrics. This suggests that moderate, relevant context benefits LLM mutation generation, whereas both insufficient and excessive context hurt efficiency and quality.
@@
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 16. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05. Although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
@@ 4.1.2. Experiment Setup of the Context Length
| Real Bug Dectc. | 85.0% | 83.3% | 83.3% |
| Avg. Ochiai | 33.3% | 35.0% | 30.0% |
| Coupling Rate | 36.6% | 35.2% | 37.1% |
| Compilability Rate | 62.9% | 63.2% | 67.2% |
| Dup. Mut. Rate | 10.8% | 12.5% | 11.9% |
@@
| GPT-3.5 P1 v.s. GPT-3.5 P2 | 0.9762 | 3.31E-05 | 0.9999 | 1.35E-31 |
| GPT-3.5 P1 v.s. GPT-3.5 P3 | 0.9040 | 2.00E-03 | 0.9999 | 8.66E-23 |
| GPT-3.5 P1 v.s. GPT-3.5 P4 | 0.9520 | 2.60E-04 | 0.9999 | 1.94E-31 |
| GPT-3.5 P1 v.s. GPT-3.5 P5 | 0.9762 | 3.31E-05 | 0.9999 | 1.57E-26 |
| DS-671b P1 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
| DS-671b P1 v.s. DS-671b P3 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
| DS-671b P1 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
| DS-671b P1 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
| 1-Line vs. 2-Line | 0.9762 | 3.31E-05 | 0.9980 | 2.12E-08 |
| 2-Line vs. 3-Line | 0.9762 | 3.31E-05 | 0.9997 | 1.04E-10 |
| 1-Line vs. 3-Line | 0.9524 | 2.60E-04 | 0.9978 | 2.69E-08 |
| QB-6a (Default) v.s. QB-6b | 0.9429 | 4.81E-03 | 0.9984 | 1.00E-08 |
| QB-6a (Default) v.s. QB-3 | 0.9856 | 3.09E-05 | 0.9999 | 1.62E-13 |
| QB-6a (Default) v.s. QB-4 | 0.9643 | 4.54E-04 | 0.9999 | 7.28E-23 |
| QB-6a (Default) v.s. QB-5 | 0.9643 | 4.54E-04 | 0.9999 | 9.62E-24 |
| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 9.39E-16 |
| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 3.34E-15 |
| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 5.01E-14 |
| QB-6a (Default) v.s. QB-10 | 0.9856 | 3.09E-04 | 0.9999 | 2.89E-13 |
| QB-6a (Default) v.s D4J-6 | 0.9524 | 2.60E-04 | 0.9980 | 2.13E-08 |
@@
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutations for all the approaches. To check the impact of context selection on performance, we conduct experiments on 60 bugs from Defects4J, comparing how different context lengths affect performance. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 15. Similarly, the middle part of Table 16 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95, and the -values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
@@ 4.1.3. Impact of Different Few-Shot Examples
| Real Bug Detec | 0.917 | 0.933 | 0.883 | 0.883 | 0.917 | 0.900 | 0.900 | 0.917 | 0.867 | 0.900 |
| Avg. Ochiai | 0.467 | 0.450 | 0.400 | 0.433 | 0.433 | 0.467 | 0.433 | 0.433 | 0.400 | 0.367 |
| Coupling Rate | 0.136 | 0.128 | 0.126 | 0.127 | 0.132 | 0.112 | 0.113 | 0.121 | 0.136 | 0.139 |
| Comp. Rate | 67.5% | 62.7% | 65.6% | 66.8% | 63.9% | 65.1% | 66.0% | 61.8% | 64.9% | 62.1% |
| Dup.Mut.Rate | 13.7% | 12.8% | 12.6% | 9.4% | 12.8% | 11.4% | 13.2% | 13.8% | 13.6% | 13.9% |
@@
To measure the impact of different few-shot examples on the results, we conduct comparative experiments with various examples. First, we randomly select 3, 4, 5, 6, 7, 8, 9, and 10 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 17.
@@
4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397. This ensures a fair comparison across approaches by controlling for differences in mutation counts, thereby eliminating the potential bias that methods generating more mutations might appear to perform better simply due to quantity rather than quality. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 18. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
| Name | R. B. Detec. | Ochiai Coeffi. | Coup. Rate | Comp. Rate | Dup. Rate |
@@ 4.1.4. Performance of Using the Same Number of Mutations
| Metric | LLMut (GPT-3.5) | LLMut (DS-761b) | ||||
| Default | More Detailed | Default + 10 Err. | Default | More Detailed | Default + 10 Err. | |
| Real Bug Detec | 0.917 | 0.900 | 0.900 | 0.950 | 0.950 | 0.933 |
| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.500 | 0.500 | 0.517 |
| Coupling Rate | 0.136 | 0.129 | 0.117 | 16.9% | 17.8% | 16.0% |
| Comp. Rate | 67.5% | 71.1% | 68.7% | 78.4% | 76.4% | 77.5% |
| Dup.Mut.Rate | 13.7% | 14.1% | 10.4% | 5.1% | 4.1% | 6.2% |
Note: The prompt achieving the best performance in terms of a metric is highlighted in grey background color.
4.2. Impact of Prompt Turbulence on LLM-Based Mutation Generation
In software engineering practice, even when using the same information sources, users may design prompts differently, for example, varying the level of detail in the requirements description or introducing minor language errors. To investigate the impact of such variations, we derived two prompt variants from the default prompt shown in Figure 1.
More Detailed Output Specification: replacing the output specifications of the default prompt by explicitly requiring: Prohibit generating identical mutants, i.e., ensure ‘aftercode’ ‘precode’, and all mutants are unique. Write all output mutants in a JSON file.
Default + 10 Errors: randomly injecting 10 English grammar mistakes into the default prompt, such as subject–verb agreement errors (e.g., it are) and noun number errors (e.g., two mutation).
Following the same experimental setup as before, we compared these two variants using GPT-3.5 and DeepSeek-V3-761b, with the results summarized in Table 19.
For GPT-3.5, the More Detailed Output Specification variant slightly improves the compilation rate (+3.6%) but also increases the duplicate mutation rate. In contrast, the Default + 10 Errors variant reduces real bug detection and coupling rate compared to the default prompt. Interestingly, the Default + 10 Errors variant, presumably the lowest in prompt quality, achieves the best performance on duplicate mutation rate. For DeepSeek-V3-761b, the differences across variants are minimal. For instance, the default and the More Detailed Output Specification prompts achieve identical real bug detection and average Ochiai, and the More Detailed Output Specification variant even shows a small increase in average Ochiai.
Overall, these results suggest that prompt wording has a certain degree of randomness in its effect on mutation generation performance. To further verify this, we conducted similarity analysis across the three prompts for each metric and found no statistically significant superiority of any one prompt over the other two.
| Metric | LLMut (GPT-3.5) | LLMut (GPT-4o) | LLMut (DS-671b) | |||
| FOM | HOM | FOM | HOM | FOM | HOM | |
| Real Bug Detec. | 0.917 | 0.900 | 0.933 | 0.933 | 0.950 | 0.933 |
| Avg. Ochiai | 0.467 | 0.483 | 0.467 | 0.483 | 0.500 | 0.483 |
| Coupling Rate | 0.136 | 0.249 | 0.151 | 0.307 | 0.169 | 0.306 |
| Compilability Rate | 67.5% | 63.4% | 76.3% | 70.2% | 78.4% | 73.3% |
| Dup. Mut. Rate | 13.7% | 5.2% | 5.3% | 3.5% | 5.1% | 3.1% |
@@ 4.2. Impact of Prompt Turbulence on LLM-Based Mutation Generation
Note: The order achieving the best performance in terms of a metric is highlighted in grey background color.
4.3. Effectiveness on High-Order Mutation Generation

### v5

@@ 4. Discussion
In this section, we examine the impact of various experiment setups for LLMs and discuss their implications.
@@ 4.1.1. Setup of the Prompts
Prompts play a critical role in the performance of LLMs. Therefore, we aim to investigate how different prompts influence mutant generation effectiveness. Due to budget constraints, we selected a subset of 306 bugs from our dataset, comprising 10 randomly sampled bugs from each Defects4J project and all bugs from the ConDefects dataset.
Based on the prompt design of LLMut, we investigate how different sources of contextual information affect mutant generation. For this experiment, we select two representative LLMs—the closed-source GPT-4o and the recent open-source DeepSeek-V3-671b—and vary the prompt template used for mutation generation. Note that based on the evaluation results, both models are leading LLMs in various metrics. Starting from the prompt that includes the richest contextual information, we progressively remove specific sources of context in the Context section. This process yields five distinct prompt variants, summarized as follows:
Prompt 1 (P1): P1 is the richest prompt with the code of the whole Java source file as the Context of the prompt.
Prompt 2 (P2): P2 is the prompt with the code of the focal method and its corresponding unit tests as the Context of the prompt.
Prompt 3 (P3): P3 is the prompt with the code of the focal method as the Context of the prompt, which is the default prompt in the previous section, shown as Figure 1.
Prompt 4 (P4): P4 is the prompt with the code of the target code element to be mutated as the Context of the prompt.
Prompt 5 (P5): P5 is the prompt with the code of the focal method as the Context, but remove all few-shot examples.
| Metric | LLMut (GPT-4o) | LLMut (DS-671b) | ||||||||
@@
| Mut. # | 3696 | 3692 | 3781 | 3807 | 3788 | 3878 | 3844 | 3851 | 3834 | 3847 |
| Mut. Score | 0.841 | 0.801 | 0.773 | 0.832 | 0.785 | 0.776 | 0.753 | 0.765 | 0.760 | 0.750 |
| Real Bug Detec. | 83.1% | 83.9% | 85.0% | 83.1% | 84.7% | 83.1% | 84.7% | 85.5% | 83.9% | 85.8% |
| Avg. Ochiai | 64.8% | 65.0% | 66.4% | 64.8% | 65.3% | 66.4% | 65.8% | 67.5% | 66.4% | 65.6% |
| Coupling Rate | 47.9% | 49.7% | 52.2% | 48.4% | 52.1% | 49.1% | 50.0% | 53.2% | 50.9% | 51.7% |
| Compilability Rate | 78.4% | 81.6% | 80.1% | 77.4% | 82.5% | 79.5% | 82.0% | 82.8% | 78.0% | 83.5% |
| Dup. Mut. Rate | 7.7% | 7.9% | 7.9% | 8.7% | 8.1% | 8.2% | 7.0% | 7.3% | 8.5% | 6.4% |
| Avg. Gen. Time | 1.607 | 1.584 | 1.389 | 1.017 | 1.282 | 2.596 | 2.601 | 2.573 | 2.488 | 2.571 |
| Avg. # Token | 392 | 218 | 105 | 44 | 94 | 392 | 219 | 104 | 43 | 92 |
@@
Note: P3 is the prompt used in the previous section. The prompt template that achieves the best performance is highlighted in grey background color.
Table 14 presents a comparative evaluation of different prompts on two models across the main effectiveness, validity, and efficiency metrics.
For the effectiveness metrics (i.e., real bug detectability, average Ochiai, and coupling rate), on both models, P3, the prompt used in the previous section, delivers the best overall performance, demonstrating its strength in generating mutants that are semantically closer to real bugs. P5, which removes the few-shot examples, ranks as the second-best prompt overall, suggesting that few-shot examples help guide the model toward generating mutants closer to real bugs. Interestingly, P1, the longest prompt with the whole Java file, performs the worst in most cases, indicating that providing more information does not necessarily help LLMs generate more bug-representative mutants. P2, which adds unit tests to the context compared with P3, shows slightly reduced effectiveness—suggesting that including unit tests does not help the model generate better mutants.
For the validity metrics, different prompts express a similar trend in both models. For the compilability rate, P5, the one without real-world bug few-shot examples, achieves the highest compilability rate on both models. This indicates that removing few-shot examples reduces the syntactic complexity of the generated mutations, making them easier for LLMs to produce without introducing compilation errors. For the duplicated mutant rate, P1 achieves the best performance on GPT-4o, while P5 performs best on DS-671b, suggesting that duplication is more influenced by model behavior than by the prompts.
For the efficiency costs, the results show a clear trend: prompts with more contextual information consume more time and tokens. For example, P1 (whole file) incurs the highest cost, followed by P3 (whole method), while P5 (only the code snippet to be mutated) is the most efficient. Another interesting finding is that P1 (whole file) is significantly longer than P2 (unit tests).
@@
We further explore the performance similarity of different prompts. To measure their similarity in performance, we use the Spearman coefficient and Pearson correlation, as shown in the top part of Table 16. We can see that their mutual similarity exceeds 0.95, which is above the typical threshold of 0.85 (Siami Namin et al., 2008; Motwani and Brun, 2023; Kaufman et al., 2022), and the -values are all below 0.05. The results suggest that on both GPT-4o and DS-671b, although P3 achieves the most balanced overall performance, the differences among prompts are not statistically significant.
@@ 4.1.2. Experiment Setup of the Context Length
| Real Bug Dectc. | 93.3% | 91.7% | 93.3% |
| Avg. Ochiai | 46.7% | 50.0% | 48.3% |
| Coupling Rate | 37.3% | 36.8% | 37.7% |
| Compilability Rate | 80.7% | 81.4% | 82.1% |
| Dup. Mut. Rate | 4.9% | 6.6% | 5.3% |
@@
| GPT-4o P3 v.s. GPT-4o P1 | 0.9500 | 8.76E-05 | 0.9969 | 5.37E-09 |
| GPT-4o P3 v.s. GPT-4o P2 | 1.0000 | 0.00E+00 | 0.9995 | 8.56E-12 |
| GPT-4o P3 v.s. GPT-4o P4 | 0.9500 | 8.76E-05 | 0.9998 | 8.32E-14 |
| GPT-4o P3 v.s. GPT-4o P5 | 1.0000 | 0.00E+00 | 0.9999 | 5.61E-19 |
| DS-671b P3 v.s. DS-671b P1 | 1.0000 | 0.00E+00 | 0.9999 | 1.51E-30 |
| DS-671b P3 v.s. DS-671b P2 | 1.0000 | 0.00E+00 | 0.9999 | 4.95E-31 |
| DS-671b P3 v.s. DS-671b P4 | 1.0000 | 0.00E+00 | 0.9999 | 2.93E-29 |
| DS-671b P3 v.s. DS-671b P5 | 1.0000 | 0.00E+00 | 0.9999 | 1.08E-31 |
| 1-Line vs. 2-Line | 0.9999 | 1.40E-24 | 0.9989 | 4.54E-05 |
| 2-Line vs. 3-Line | 0.9999 | 1.40E-24 | 0.9995 | 1.46E-05 |
| 1-Line vs. 3-Line | 0.9999 | 1.40E-24 | 0.9998 | 2.99E-06 |
| QB-6a (Default) v.s. QB-6b | 1.0000 | 0.00E+00 | 0.9999 | 2.99E-15 |
| QB-6a (Default) v.s. QB-3 | 1.0000 | 0.00E+00 | 0.9999 | 1.82E-14 |
| QB-6a (Default) v.s. QB-4 | 1.0000 | 0.00E+00 | 0.9999 | 3.97E-15 |
| QB-6a (Default) v.s. QB-5 | 1.0000 | 0.00E+00 | 0.9999 | 7.62E-14 |
| QB-6a (Default) v.s. QB-7 | 1.0000 | 0.0E+00 | 0.9999 | 1.12E-14 |
| QB-6a (Default) v.s. QB-8 | 1.0000 | 0.0E+00 | 0.9999 | 1.28E-13 |
| QB-6a (Default) v.s. QB-9 | 1.0000 | 0.00E+00 | 0.9999 | 7.03E-15 |
| QB-6a (Default) v.s. QB-10 | 1.0000 | 0.00E+00 | 0.9999 | 7.03E-15 |
| QB-6a (Default) v.s D4J-6 | 1.0000 | 0.00E+00 | 0.9999 | 9.81E-17 |
@@
To compare the similarity to real defects, we deliberately select the context around the bug locations for generating mutants for all the approaches. In our previous experiments, the context length is set to three lines around the bug location, so we conduct experiments to compare the effects of two-line and one-line contexts, as shown in Table 15. Similarly, the middle part of Table 16 shows the performance similarity among different contexts. We can see that their mutual similarity exceeds 0.95, and the -values are all below 0.05. Different context lengths perform similarly, thus validating the setup of our experiment.
@@ 4.1.3. Impact of Different Few-Shot Examples
| Real Bug Detec | 0.933 | 0.900 | 0.917 | 0.933 | 0.917 | 0.900 | 0.883 | 0.933 | 0.933 | 0.933 |
| Avg. Ochiai | 0.483 | 0.483 | 0.450 | 0.417 | 0.450 | 0.450 | 0.433 | 0.467 | 0.483 | 0.483 |
| Coupling Rate | 0.379 | 0.379 | 0.385 | 0.380 | 0.373 | 0.362 | 0.348 | 0.367 | 0.352 | 0.388 |
| Comp. Rate | 82.1% | 82.4% | 80.8% | 81.7% | 81.6% | 81.4% | 81.8% | 82.2% | 81.4% | 83.1% |
| Dup.Mut.Rate | 5.3% | 5.6% | 6.4% | 5.6% | 5.8% | 5.6% | 7.5% | 6.4% | 6.2% | 5.8% |
@@
To evaluate how different few-shot examples influence the results, we conduct comparative experiments by varying the few-shot examples used in the default LLMut (GPT-4o) prompt. First, we randomly select 3, 4, 5, 6, 7, 8, 9, and 10 additional examples from QuixBugs, and ensure there is no overlap with the default 6 examples. Additionally, despite the risk of data leakage, we select 6 examples from the Defects4J dataset. The results are presented in Table 17.
@@
4.1.4. Performance of Using the Same Number of Mutants
In Section 3, we did not restrict the number of mutants generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of the existing studies (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutants generated by all methods to the minimum produced by BERT (Degiovanni and Papadakis, 2022), which is 11397. This ensures a fair comparison across approaches by controlling for differences in mutant counts, thereby eliminating the potential bias that methods generating more mutants might appear to perform better simply due to quantity rather than quality. For counts exceeding this number, we randomly selected the specified number of mutants for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 18.
Under a fixed number of mutants, we observe that LLMut (DS-671b) continues to deliver the best performance in both real bug detection rate and average Ochiai coefficient, with LLMut (GPT-4o) following closely. LLMorpheus (DS-671b) and LLMorpheus (GPT-4o) achieve the highest coupling rates. In terms of validity metrics—compilability rate and duplicate mutant rate—BugFarm (DS-671b) performs the best among all LLM-based approaches. These findings indicate that, when controlling for the number of generated mutants, different LLM-based approaches (e.g., LLMut, LLMorpheus, and BugFarm) exhibit distinct strengths rather than a single universally superior approach.
| Name | R. B. Detec. | Avg. Ochiai | Coup. Rate | Comp. Rate | Dup. Rate |
@@ 4.1.4. Performance of Using the Same Number of Mutants
| Metric | LLMut (GPT-4o) | LLMut (DS-671b) | ||
| FOM | HOM | FOM | HOM | |
| Real Bug Detec. | 0.933 | 0.933 | 0.950 | 0.933 |
| Avg. Ochiai | 0.483 | 0.483 | 0.500 | 0.483 |
| Coupling Rate | 0.379 | 0.507 | 0.377 | 0.506 |
| Compilability Rate | 82.1% | 70.2% | 84.1% | 73.3% |
| Dup. Mut. Rate | 5.3% | 3.5% | 5.1% | 3.1% |
@@
Note: The order achieving the best performance in terms of a metric is highlighted in a grey background color.
4.2. Effectiveness on High-Order Mutant Generation